Safawi kicked a ball with an initial speed of 20 ms-1 at 30º angle. If the ball experienced a constant vertical acceleration of -9.81 ms-2, calculate the ball’s maximum height and distance. Explain how soccer player could apply the force of summation theory to optimize his kicks performance.



1
Expert's answer
2022-02-08T10:02:16-0500

"h=\frac{v^2\sin^2\alpha}{2g}=5.1~m,"

"l=\frac{v^2\sin2\alpha}g=35.3~m."